OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y = x +

OpenStudy (anonymous):

@princeharryyy

OpenStudy (princeharryyy):

y=2x+c 4= 2*3 + c c=-2 y=2x-2

OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (3, 1) and (2, 9)? y = x +

OpenStudy (princeharryyy):

y = -8x +c 1 = -24 +c y=-8x + 25

OpenStudy (princeharryyy):

i gotta go to sleep it's 5 am, talk u later .